# Asynchronous Federated Reinforcement Learning with Policy Gradient Updates: Algorithm Design and Convergence Analysis

### v1

@@ Abstract.
To improve the efficiency of reinforcement learning, we propose a novel asynchronous federated reinforcement learning framework termed AFedPG, which constructs a global model through collaboration among $N$ agents using policy gradient (PG) updates. To handle the challenge of lagged policies in asynchronous settings, we design delay-adaptive lookahead and normalized update techniques that can effectively handle the heterogeneous arrival times of policy gradients. We analyze the theoretical global convergence bound of AFedPG, and characterize the advantage of the proposed algorithm in terms of both the sample complexity and time complexity. Specifically, our AFedPG method achieves $\mathcal{O}(\frac{{\epsilon}^{-2.5}}{N})$ sample complexity at each agent on average. Compared to the single agent setting with $\mathcal{O}(\epsilon^{-2.5})$ sample complexity, it enjoys a linear speedup with respect to the number of agents. Moreover, compared to synchronous FedPG, AFedPG improves the time complexity from $\mathcal{O}(\frac{t_{\max}}{N})$ to $\mathcal{O}(\frac{1}{\sum_{i=1}^{N}\frac{1}{t_{i}}})$, where $t_{i}$ denotes the time consumption in each iteration at the agent $i$, and $t_{\max}$ is the largest one. The latter complexity $\mathcal{O}(\frac{1}{\sum_{i=1}^{N}\frac{1}{t_{i}}})$ is always smaller than the former one, and this improvement becomes significant in large-scale federated settings with heterogeneous computing powers ($t_{\max}\gg t_{\min}$). Finally, we empirically verify the improved performances of AFedPG in three MuJoCo environments with varying numbers of agents. We also demonstrate the improvements with different computing heterogeneity.
Reinforcement Learning, Policy Gradient, Federated Learning, Asynchronous, Optimization
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; 2024; isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learningccs: Computing methodologies Artificial intelligence Distributed artificial intelligenceccs: Computer systems organization Distributed architecturesccs: Networks Network services
1. Introduction
@@
Policy gradient (PG) methods, also known as REINFORCE (Williams, 1992), are widely used to solve reinforcement learning (RL) problems across a range of applications, with recent notable examples including reinforcement learning from human feedback (RLHF) in Google’s Gemma (Team and DeepMind, 2024), OpenAI’s InstructGPT (Ouyang et al., 2022), and ChatGPT (GPT-4 (OpenAI, 2023)). It has also garnered significant attention in communications/networking communities for solving network optimization (Wei et al., 2023; Hu et al., 2022) and wireless communication (Yu et al., 2023; Xiong et al., 2022) problems.
@@
Most practical RL applications are on large scales and rely on a huge amount of data samples for model training in an online behavior (Provost and Fawcett, 2013; Liu et al., 2021), which is considered as the key bottleneck of RL (Dulac-Arnold et al., 2021; Ladosz et al., 2022). To reduce sample complexity while enabling training on massive data, a conventional approach involves transmitting locally collected samples from distributed agents to a central server (Predd et al., 2006). The transmitted samples can then be used for policy learning on the server side. However, this may not be feasible in real-world mobile systems where the communication bandwidth is limited and long delays are intolerable (Hu et al., 2024; Lan et al., 2023a; Niknam et al., 2020), such as wireless edge devices (Han et al., 2023; Lan et al., 2023a), Internet of Things (Nguyen et al., 2021; Fang et al., 2023), autonomous driving (Kiran et al., 2022), and vehicle transportation (Al-Abbasi et al., 2019; Haliem et al., 2021). In RL, since new data is continuously generated based on the current policy, the samples collected at the agents need to be transmitted frequently to the server throughout the training process, resulting in a significant bottleneck with long delays. Furthermore, the sharing of individual data collected by agents may lead to privacy and legal issues (Kairouz et al., 2021; Mothukuri et al., 2021).
@@
Federated learning (FL) (Hosseinalipour et al., 2020; McMahan et al., 2017) offers a promising solution to the above challenges in a distributed setting. In FL, instead of directly transmitting the raw datasets to the central server, agents only communicate locally trained model parameters (or gradients) with the server. Although FL has been mostly studied for supervised learning problems, several recent works expanded the scope of FL to federated reinforcement learning (FedRL), where $N$ agents collaboratively learn a global policy without sharing the trajectories they collected during agent-environment interaction (Jin et al., 2022; Khodadadian et al., 2022; Xie and Song, 2023). FedRL has been studied in the tabular case (Agarwal et al., 2021a; Jin et al., 2022; Khodadadian et al., 2022), and for value-function based algorithms (Wang et al., 2023; Xie and Song, 2023), where a linear speedup has been demonstrated. For policy-gradient based algorithms, namely FedPG, we note that linear speedup is easy to see as the trajectories collected by each agent could be parallelized (Lan et al., 2023b). However, policy-gradient based FedRL has not been well studied in terms of global (non-local) convergence and time complexity.
@@
Despite all the aforementioned works in FedRL, they still face challenges in terms of time complexity, primarily due to their focus on synchronous model aggregation. In large-scale heterogeneous settings (Xie et al., 2019; Chen et al., 2020), performing synchronous global updates has limitations, as the overall time consumption heavily depends on the slow agents, i.e., stragglers (Mishchenko et al., 2022). In this paper, we aim to tackle this issue by strategically leveraging asynchronous federated learning (A-FL) in policy-based FedRL for the first time. A-FL (Xie et al., 2019) shows superiority compared to synchronous FL, and recent works (Mishchenko et al., 2022; Koloskova et al., 2022) further improve convergence performances with theoretical guarantees. However, compared to prior A-FL approaches focusing on supervised learning, integrating A-FL with policy-based FedRL introduces new challenges due to the presence of lagged policies in asynchronous settings. Unlike supervised FL where the datasets of the clients are fixed, in RL, agents collect new samples in each iteration based on the current policy. This dynamic nature of the data collection process makes both the problem itself and the theoretical analysis challenging. To the best of our knowledge, this problem setting and its challenges have been largely overlooked in existing research, despite the significance of employing FL in RL. The key question that this paper aims to address is:
@@
Table 1. Performance improvements of our AFedPG over other policy gradient methods. We compare the complexity for global convergence in each agent.
PG FedPG AFedPG (ours)
Sample Complexity $\mathcal{O}({\epsilon}^{-2.5})$ $\mathcal{O}(\frac{{\epsilon}^{-2.5}}{N})$ $\mathcal{O}(\frac{{\epsilon}^{-2.5}}{N})$
Global Time - $\mathcal{O}(\frac{t_{\max}}{N}{\epsilon}^{-2.5})$ $\mathcal{O}(\bar{t}{\epsilon}^{-2.5})$
@@
We answer this question in the affirmative by proposing AFedPG, an algorithm that asynchronously updates the global policy using policy gradients from federated agents. The key components of the proposed AFedPG include a delay-adaptive lookahead technique, which addresses the inconsistent arrival times of updates during the training process. The improvement of AFedPG over other approaches is summarized in Table 1. Here, the global time is measured by the number of global updates $\times$ time complexity in each iteration. In terms of sample complexity, the federated learning technique brings a linear speedup with respect to the number of agents $N$. As for the global time, AFedPG improves from $\mathcal{O}(\frac{t_{\max}}{N}{\epsilon}^{-2.5})$ to $\mathcal{O}(\bar{t}{\epsilon}^{-2.5})$, where $\bar{t}\coloneqq 1/\sum_{i=1}^{N}\frac{1}{t_{i}}$ is a harmonic average and always smaller than $t_{\max}/N$, and $t_{i}$ denotes the time complexity in each iteration at agent $i$.
1.1. Summary of Contributions
@@
  1. (1)
    New methodology with a delay-adaptive technique: We propose AFedPG, an asynchronous training method tailored to FedRL. To handle the delay issue in the asynchronous FedRL setting, we design a delay-adaptive lookahead technique. Specifically, in the $k$-th iteration of training, the agent collects samples according to the local model parameters $\widetilde{\theta}_{k}\leftarrow\theta_{k}+\frac{1-\alpha_{k-\delta_{k}}}{\alpha_{k-\delta_{k}}}(\theta_{k}-\theta_{k-1})$, where $\delta_{k}$ is the delay. This updating technique cancels out the second-order correction terms and thus assists the convergence analysis.
  2. (2)
    Convergence analysis: This work gives the global convergence guarantees of the federated policy-based RL for the first time. We analytically characterize the convergence bound of AFedPG using the key lemmas, and show the impact of various parameters including delay and number of iterations in our analysis.
  3. (3)
    Linear speedup in sample complexity: As shown in Table 1, our federated method improves the sample complexity in each agent from $\mathcal{O}({\epsilon}^{-2.5})$ (single agent PG) to $\mathcal{O}(\frac{{\epsilon}^{-2.5}}{N})$, where $N$ is the number of federated agents. This represents the linear speedup with respect to the number of agents $N$.
  4. (4)
    Time complexity improvement: Our AFedPG also reduces the time complexity of synchronous FedPG from $\mathcal{O}(\frac{t_{\max}}{N})$ to $\mathcal{O}(\bar{t}\coloneqq\frac{1}{\sum_{i=1}^{N}\frac{1}{t_{i}}})$. The latter is always smaller than the former. This gap is significant in large-scale federated settings with heterogeneous computing powers ($t_{\max}\gg t_{\min}$).
  5. (5)
    Experiments under MuJoCo environment: We empirically verify the improved performances of AFedPG in three different MuJoCo environments with varying numbers of agents. We also demonstrate the improvements with different computing heterogeneity.
To the best of our knowledge, this is the first work to successfully integrate policy-based reinforcement learning with asynchronous federated learning and analyze its behavior. This new setting necessitates us to deal with the lagged policies under a time-varying data scenario depending on the current policy.
@@
The remainder of this paper is organized as follows. In the next subsection, we summarize the related works. In Section 2, we introduce our problem setting, including some backgrounds on reinforcement learning and policy gradients. In Section 3, we describe our proposed AFedPG method. In Section 4, we analyze the global convergence rate of AFedPG and discuss its time complexity compared to the synchronous setting. Section 5 presents the experimental results, and we conclude this paper in Section 6.
1.2. Related Work
@@
In this subsection, we review previous works that are most relevant to the present work.
@@
Policy gradient methods: For vanilla PG, the state-of-the-art result is presented in (Yuan et al., 2022), achieving a sample complexity of $\widetilde{\mathcal{O}}({\epsilon}^{-4})$ for the local convergence. Several recent works have improved this boundary with PG variants. In (Huang et al., 2020), a PG with momentum method is proposed with convergence rate $\mathcal{O}({\epsilon}^{-3})$ for the local convergence. The authors of (Ding et al., 2022) further improve the convergence analysis of PG with momentum and achieve a global convergence with the rate $\mathcal{O}({\epsilon}^{-3})$. In (Fatkhullin et al., 2023), a normalized PG technique is introduced and improves the global convergence rate to $\mathcal{O}({\epsilon}^{-2.5})$. In (Mondal and Aggarwal, 2024), an acceleration-based natural policy gradient method is proposed with sample complexity $\mathcal{O}({\epsilon}^{-2})$, but second-order matrices are computed in each iteration (with first-order information), which brings more computational cost. However, all previous works have focused on the single-agent scenario, and the federated PG has not been explored. Compared to these works, we focus on a federated PG setting to improve the efficiency of RL in practical settings with distributed agents. Our scheme achieves linear speedup with respect to the number of agents, significantly reducing the sample complexity compared to the conventional PG approaches.
@@
Asynchronous FL: In (Xie et al., 2019), the superiority of asynchronous federated learning (A-FL) has been empirically shown compared to synchronous FL in terms of convergence performances. In (Chen et al., 2020), the asynchronous analysis is extended to the vertical FL with a better convergence performance compared to the synchronous setting. In (Dun et al., 2023), a dropout regularization method is introduced to handle the heterogeneous problems in A-FL. About the same time, recent works (Mishchenko et al., 2022; Koloskova et al., 2022) further improve the convergence performance with theoretical guarantees, and show that A-FL always exceeds synchronous FL without any changes to the algorithm. We note that all previous A-FL works focus on supervised learning with fixed datasets on the client side. However, in RL, agents collect new samples that depend on the current policy (model parameters) in each iteration, which makes the problem fundamentally different and challenging. In this work, we address this challenge by developing an A-FL method highly tailored to policy gradient.
@@
FedRL: For value-function based algorithms, (Jin et al., 2022) analyzes the convergence performances with environment heterogeneity. In (Agarwal et al., 2021a), a local update method is used to reduce communication costs. (Khodadadian et al., 2022) shows a linear speedup with respect to the number of agents. All of the above works are limited to tabular RL analysis. For actor-critic based method, (Wang et al., 2023) analyzes the convergence performances with linear function approximation. (Mnih et al., 2016) builds a practical system to implement the neural network approximation. In (Xie and Song, 2023), it adds KL divergence and experimentally validates the actor-critic based method with neural network approximation. For policy-based methods, in (Lan et al., 2023b), it shows simplicity compared to the other RL methods, and a linear speedup has been demonstrated in the synchronous setting. However, with an online behavior, it is not practical to perform synchronous global updates with heterogeneous computing power, and the global time consumption heavily depends on the stragglers (Mishchenko et al., 2022). This motivates us to consider the asynchronous policy-based method for FedRL. We show that our method can further reduce the time complexity compared to the synchronous FedRL approach.
2. Problem Setup

### v5

@@ Abstract
To improve the efficiency of reinforcement learning (RL), we propose a novel asynchronous federated reinforcement learning (FedRL) framework termed AFedPG, which constructs a global model through collaboration among $N$ agents using policy gradient (PG) updates. To address the challenge of lagged policies in asynchronous settings, we design a delay-adaptive lookahead technique specifically for FedRL that can effectively handle heterogeneous arrival times of policy gradients. We analyze the theoretical global convergence bound of AFedPG, and characterize the advantage of the proposed algorithm in terms of both the sample complexity and time complexity. Specifically, our AFedPG method achieves $\mathcal{O}(\frac{{\epsilon}^{-2.5}}{N})$ sample complexity for global convergence at each agent on average. Compared to the single agent setting with $\mathcal{O}(\epsilon^{-2.5})$ sample complexity, it enjoys a linear speedup with respect to the number of agents. Moreover, compared to synchronous FedPG, AFedPG improves the time complexity from $\mathcal{O}(\frac{t_{\max}}{N})$ to $\mathcal{O}({\sum_{i=1}^{N}\frac{1}{t_{i}}})^{-1}$, where $t_{i}$ denotes the time consumption in each iteration at agent $i$, and $t_{\max}$ is the largest one. The latter complexity $\mathcal{O}({\sum_{i=1}^{N}\frac{1}{t_{i}}})^{-1}$ is always smaller than the former one, and this improvement becomes significant in large-scale federated settings with heterogeneous computing powers ($t_{\max}\gg t_{\min}$). Finally, we empirically verify the improved performance of AFedPG in four widely used MuJoCo environments with varying numbers of agents. We also demonstrate the advantages of AFedPG in various computing heterogeneity scenarios.
1 Introduction
@@
Policy gradient (PG) methods, also known as REINFORCE (Williams, 1992), are widely used to solve reinforcement learning (RL) problems across a range of applications, with recent notable examples including reinforcement learning from human feedback (RLHF) in Google’s Gemma (Team & DeepMind, 2024), OpenAI’s InstructGPT (Ouyang et al., 2022), and ChatGPT (GPT-4 (OpenAI, 2023)). PG-based methods has also garnered significant attention in several applications including transportation systems Al-Abbasi et al. (2019), networking Geng et al. (2023), data-center resource allocation Chen et al. (2023; 2024a); Zhang et al. (2023), streaming Elgabli et al. (2024); Liu & Zhu (2023), robotics Gonzalez et al. (2023), and diffusion models Zhang* et al. (2024).
@@
Most practical RL applications operate at large scales and rely on a huge amount of data samples for model training in an online behavior (Provost & Fawcett, 2013; Liu et al., 2021), which is considered as a key bottleneck in RL (Dulac-Arnold et al., 2021; Ladosz et al., 2022). To reduce sample complexity while enabling training on massive data, a conventional approach involves transmitting locally collected samples from distributed agents to a central server (Predd et al., 2006). The transmitted samples can then be used for policy learning on the server side. However, this may not be feasible in real-world mobile systems where the communication bandwidth is limited and a large training time is intolerable (Brinton et al., 2024; Chu et al., 2024a; Yuan et al., 2024; Niknam et al., 2020; Elgabli et al., 2020), such as wireless edge devices (Lan et al., 2023a; Ching et al., 2023; 2024), cloud computing (Chen et al., 2024b; Xu et al., 2024), autonomous driving (Kiran et al., 2022), and financial applications (Long et al., 2024; Liang et al., 2024). In RL, since new data is continuously generated based on the current policy, the samples collected at the agents need to be transmitted frequently to the server throughout the training process (Shen et al., 2023), resulting in a significant bottleneck with large delays. Furthermore, the sharing of individual data collected by agents may lead to privacy and legal issues (Kairouz et al., 2021; Mothukuri et al., 2021).
@@
Federated learning (FL) (McMahan et al., 2017) offers a promising solution to the above challenges in a distributed setting. In FL, instead of directly transmitting the raw datasets to the central server, agents only communicate locally trained model parameters (or gradients) with the server. In Chu et al. (2025; 2024b), it shows that FL could achieve fairness among different agents. Although FL has been mostly studied for supervised learning problems, several recent works expanded the scope of FL to federated reinforcement learning (FedRL), where $N$ agents collaboratively learn a global policy without sharing the trajectories they collected during agent-environment interaction (Jin et al., 2022; Khodadadian et al., 2022; Xie & Song, 2023). FedRL has been studied in the tabular case (Agarwal et al., 2021b; Jin et al., 2022; Khodadadian et al., 2022), and for value-function based algorithms (Wang et al., 2024b; Xie & Song, 2023), where a linear speedup has been demonstrated. For policy-gradient based algorithms, namely FedPG, we note that a linear speedup is easy to achieve, given that trajectories collected at different agents can be processed in parallel (Lan et al., 2023b). Ganesh et al. (2024a) proposed a policy gradient-based approach that is robust to adversarial agents and achieves optimal sample complexity in the presence of adversaries. Liu & Zhu (2022) analyzes the performances in the multi-agent setting. Zhu et al. (2024) extends the result to the multi-task setting. However, policy-gradient based FedRL has not been well studied in terms of global (non-local) time complexity.
@@
Despite all the aforementioned works in FedRL, they still face challenges in terms of time complexity, primarily due to their focus on synchronous model aggregation. In large-scale heterogeneous settings (Xiong et al., 2024; Xie et al., 2020; Chen et al., 2020), performing synchronous global updates has limitations, as the overall time consumption heavily depends on the slow agents, i.e., stragglers (Badita et al., 2021; Mishchenko et al., 2022). In this paper, we aim to tackle this issue by strategically leveraging asynchronous federated learning (A-FL) in policy-based FedRL for the first time. A-FL (Xie et al., 2020; Chang et al., 2024) shows superiority compared to synchronous FL, and recent works (Mishchenko et al., 2022; Koloskova et al., 2022) further improve convergence performances with theoretical guarantees.
Challenges: However, compared to prior A-FL approaches focusing on supervised learning, integrating A-FL with policy-based FedRL introduces new challenges due to the presence of lagged policies in asynchronous settings. Unlike supervised FL where the datasets of the clients are fixed, in RL, agents collect new samples in each iteration based on the current policy. This dynamic nature of the data collection process makes both the problem itself and the theoretical analysis challenging. Guaranteeing the global convergence of the algorithm is especially non-trivial under the asynchronous FedRL setting with lagged policies. This problem setting and its challenges have been largely overlooked in existing research, despite the significance of employing FL in RL. The key question that this paper aims to address is:
@@
Table 1: Performance improvements of our AFedPG over other first-order policy gradient methods. We compare the complexity for convergence in each agent.
Sample Complexity Sample Complexity
Methods (FOSP) (Global) Global Time
Vanilla PG
(Yuan et al., 2022) $\mathcal{O}({\epsilon}^{-4})$ $\mathcal{O}({\epsilon}^{-3})$
Normalized PG
(Fatkhullin et al., 2023) $\mathcal{O}({\epsilon}^{-3.5})$ $\mathcal{O}({\epsilon}^{-2.5})$
FedPG $\mathcal{O}(\frac{{\epsilon}^{-3.5}}{N})$ $\mathcal{O}(\frac{{\epsilon}^{-2.5}}{N})$ $\mathcal{O}(\frac{t_{\max}}{N}{\epsilon}^{-2.5})$
AFedPG $\mathcal{O}(\frac{{\epsilon}^{-3.5}}{N})$ $\mathcal{O}(\frac{{\epsilon}^{-2.5}}{N})$ $\mathcal{O}(\bar{t}{\epsilon}^{-2.5})$
@@
We answer this question in the affirmative by proposing AFedPG, an algorithm that asynchronously updates the global policy using policy gradients from federated agents. The key components of the proposed approach include a delay-adaptive lookahead technique tailored to PG, which addresses the inconsistent arrival times of updates during the training process. This new approach eliminates the second-order correction terms that do not appear in conventional supervised FL, effectively addressing the unique challenges of asynchronous FedRL. The improvement of AFedPG over other approaches is summarized in Table 1. Here, the global time is measured by the number of global updates $\times$ time complexity in each iteration. In terms of sample complexity, the federated learning technique brings a linear speedup with respect to the number of agents $N$. As for the global time, AFedPG improves from $\mathcal{O}(\frac{t_{\max}}{N}{\epsilon}^{-2.5})$ to $\mathcal{O}(\bar{t}{\epsilon}^{-2.5})$, where $\bar{t}\coloneqq 1/\sum_{i=1}^{N}\frac{1}{t_{i}}$ is a harmonic average which is less than or equal to $t_{\max}/N$, and $t_{i}$ denotes the time complexity in each iteration at agent $i$.
1.1 Summary of Contributions
@@
  1. 1.
    New methodology with a delay-adaptive technique: We propose AFedPG, an asynchronous training method tailored to FedRL. To handle the delay issue in the asynchronous FedRL setting, we design a delay-adaptive lookahead technique. Specifically, in the $k$-th iteration of training, the agent collects samples according to the local model parameters $\widetilde{\theta}_{k}\leftarrow\theta_{k}+\frac{1-\alpha_{k-\delta_{k}}}{\alpha_{k-\delta_{k}}}(\theta_{k}-\theta_{k-1})$, where $\delta_{k}$ is the delay. Unlike (supervised) FL, a second-order correction term (marked blue in equation 32) only occurs in RL because of the sampling mechanism. This updating technique cancels out the second-order correction terms ($(1-\alpha_{k-\delta_{k}})\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k})+\alpha_{k-\delta_{k}}\nabla^{2}J(\theta_{k})(\widetilde{\theta}_{k}-\theta_{k})=0$) and thus assists the convergence analysis. This technique is specifically designed for AFedRL, and not developed by previous FL works.
  2. 2.
    Convergence analysis: This work gives both the global and the first-order stationary point (FOSP) convergence guarantees of the asynchronous federated policy-based RL for the first time. We analytically characterize the convergence bound of AFedPG using the key lemmas, and show the impact of various parameters including delay and number of iterations.
  3. 3.
    Linear speedup in sample complexity: As shown in Table 1, our AFedPG approach improves the sample complexity in each agent from $\mathcal{O}({\epsilon}^{-2.5})$ (single agent PG) to $\mathcal{O}(\frac{{\epsilon}^{-2.5}}{N})$, where $N$ is the number of federated agents. This represents the linear speedup of our method with respect to the number of agents $N$.
  4. 4.
    Time complexity improvement: Our AFedPG also reduces the time complexity of synchronous FedPG from $\mathcal{O}(\frac{t_{\max}}{N})$ to $\mathcal{O}(\bar{t}\coloneqq\frac{1}{\sum_{i=1}^{N}\frac{1}{t_{i}}})$. The latter is always smaller than the former. This improvement is significant in large-scale federated settings with heterogeneous delays ($t_{\max}\gg t_{\min}$).
  5. 5.
    Experiments under the MuJoCo environment: We empirically verify the improved performances of AFedPG in four different MuJoCo environments with varying numbers of agents. We also demonstrate the improvements with different computing heterogeneity.
To the best of our knowledge, this is the first work to successfully integrate policy-based reinforcement learning with asynchronous federated learning and analyze its behavior, accompanied by theoretical convergence guarantees. This new setting necessitates us to deal with the lagged policies under a time-varying data scenario depending on the updated policy.
@@
2 Related Work
@@
In this section, we review previous works that are most relevant to the present work.
@@
Policy gradient methods: For vanilla PG, the state-of-the-art result is presented in Yuan et al. (2022), achieving a sample complexity of $\widetilde{\mathcal{O}}({\epsilon}^{-4})$ for the local convergence. Several recent works have improved this boundary with PG variants. In Huang et al. (2020), a PG with momentum method is proposed with convergence rate $\mathcal{O}({\epsilon}^{-3})$ for the local convergence. The authors of Ding et al. (2022) further improve the convergence analysis of PG with momentum and achieve a global convergence with the rate $\mathcal{O}({\epsilon}^{-3})$. In Fatkhullin et al. (2023), a normalized PG technique is introduced and improves the global convergence rate to $\mathcal{O}({\epsilon}^{-2.5})$. In Mondal & Aggarwal (2024), an acceleration-based natural policy gradient method is proposed with sample complexity $\mathcal{O}({\epsilon}^{-2})$, but second-order matrices are computed in each iteration (with first-order information), which brings more computational cost. However, all previous works have focused on the single-agent scenario, and the federated PG has not been explored. Compared to these works, we focus on a practical federated PG setting with distributed agents to improve the efficiency of RL. Our scheme achieves a linear speedup with respect to the number of agents, significantly reducing the sample complexity compared to the conventional PG approaches.
@@
Asynchronous FL: In Xie et al. (2020), the superiority of asynchronous federated learning (A-FL) has been empirically shown compared to synchronous FL in terms of convergence performances. In Chen et al. (2020), the asynchronous analysis is extended to the vertical FL with a better convergence performance compared to the synchronous setting. In Dun et al. (2023), a dropout regularization method is introduced to handle the heterogeneous problems in A-FL. About the same time, recent works (Mishchenko et al., 2022; Koloskova et al., 2022) further improve the convergence performance with theoretical guarantees, and show that A-FL always exceeds synchronous FL without any changes to the algorithm. We note that all previous A-FL works focus on supervised learning with fixed datasets on the client side. However, in RL, agents collect new samples that depend on the current policy (model parameters) in each iteration, which makes the problem fundamentally different and challenging. In this work, we address this challenge by developing an A-FL method highly tailored to policy gradient, leveraging the proposed delay-adaptive lookahead technique.
@@
FedRL: For value-function based algorithms, Zheng et al. (2024); Jin et al. (2022); Woo et al. (2023) analyze the convergence performances with environment heterogeneity. Salgia & Chi (2024) analyzes the trade-off between sample and communication complexity. Zheng et al. (2024); Khodadadian et al. (2022) shows a linear speedup with respect to the number of agents. Zhang et al. (2024) extends the result with linear function approximation. However, all of the above works are limited to tabular or linear approximation analysis (without deep learning). For actor-critic (AC) based method, Wang et al. (2024b) analyzes the convergence performances with the linear function approximation. Mnih et al. (2016) builds a practical system to implement the neural network approximation. Yang et al. (2024) analyzes the sample complexity in the multi-task setting. In Xie & Song (2023), it adds KL divergence and experimentally validates the actor-critic based method with neural network approximation. For policy-based methods, Chen et al. (2021) gives a convergence guarantee for the vanilla FedPG. Wang et al. (2024a) analyzes the performance in the heterogeneous setting. Lan et al. (2023b) further shows the simplicity compared to the other RL methods, and a linear speedup has been demonstrated in the synchronous setting. Further, optimal sample complexity for global optimality in federated RL even in the presence of adversaries is studied in Ganesh et al. (2024a). However, with online behavior, it is not practical to perform synchronous global updates with heterogeneous computing power, and the global time consumption heavily depends on the stragglers (Mishchenko et al., 2022). This motivates us to consider the asynchronous policy-based method for FedRL. We demonstrate both theoretically and empirically that our method further reduces the time complexity compared to the synchronous FedRL approach.
3 Problem Setup